# SCoTER: Structured Chain-of-Thought Transfer for Enhanced Recommendation

### v4

@@ 4.2.1. Structured Distillation
This stage aims to preserve step-wise structural information and thereby retaining the information captured by $I(C;Y|S,P)$. We achieve this through a structured teacher-student distillation framework.
We leverage the optimal template to guide a powerful teacher LLM (e.g., DeepSeek-R1) in generating structured reasoning chains. For each user sequence $S$ in our training corpus, the teacher model produces template-consistent reasoning $C=(c_{1},c_{2},\ldots,c_{K})$, creating training pairs $\{(S_{i},C_{i})\}_{i=1}^{N}$ where the student learns to generate structured reasoning given user sequences as input.
A smaller, more efficient student model (Qwen3-8B (Yang et al., 2025a)) is fine-tuned on this synthetic dataset, enabling it to generate pattern-consistent reasoning chains that adapt to specific user contexts.
@@
For each generated reasoning step $c_{i,j}$, we extract a dense embedding using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025b)). This process transforms textual reasoning steps into fixed-dimensional embeddings, where $\mathbf{e}_{i,j}\in\mathbb{R}^{D}$ represents the embedding for the $j$-th reasoning step of sequence $S_{i}$:
@@
All structured embedding matrices $\{\mathbf{H}_{i}\}_{i=1}^{N}$ are computed and stored offline, enabling the lightweight online fusion phase. This allows for rapid retrieval and integration of pre-computed reasoning representations without incurring generation latency.
@@ 4.2.2. Order-Preserving Fusion
This stage integrates pre-computed step-wise embeddings with backbone recommendation models using a lightweight, model-agnostic fusion architecture. This online component prioritizes serving efficiency while preserving the sequential structure critical for reasoning effectiveness.
@@
To preserve the sequential dependencies critical for structured reasoning, we augment each projected embedding with learnable positional encodings:
@@
where $\mathbf{P}j\in\mathbb{R}^{d{\text{item}}}$ are position embeddings that encode each step’s role within the reasoning sequence.
We employ cross-attention to allow each sequence position to selectively attend to relevant reasoning steps. Let $\mathbf{e}_{\text{seq}}=[\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{T}]$ $\in\mathbb{R}^{T\times d_{\text{item}}}$ denote the backbone model’s embeddings for the user sequence, and $\mathbf{Z}^{\text{pos}}=[\mathbf{z}_{1}^{\text{pos}},\ldots,\mathbf{z}_{K}^{\text{pos}}]\in\mathbb{R}^{K\times d_{\text{item}}}$ represent the projected CoT embeddings with positional encoding. In this cross-attention, the sequence embeddings serve as queries, while the reasoning steps act as both keys and values. The cross-attention mechanism computes attended reasoning representations for each sequence position:
(5) $\mathbf{A}=\text{Attention}(\mathbf{e}_{\text{seq}},\mathbf{Z}^{\text{pos}},\mathbf{Z}^{\text{pos}})$
@@
(6) $\mathbf{g}$ $=\sigma(\mathbf{W}_{g}[\mathbf{e}_{\text{seq}};\mathbf{A}]+\mathbf{b}_{g})$
(7) $\mathbf{E}_{\text{fused}}$ $=\text{LayerNorm}(\mathbf{g}\odot\mathbf{e}_{\text{seq}}+(1-\mathbf{g})\odot\mathbf{A})$
Here, $\mathbf{e}_{\text{seq}};\mathbf{A}]$ represents concatenation along the feature dimension, and the final layer normalization is applied to the gated output.
To align the reasoning space with the recommendation objective, we employ a contrastive learning component using the InfoNCE loss (Oord et al., 2018). The loss is computed between the final reasoning step embedding, $\mathbf{z}_{K}$, and the target item embedding, $\mathbf{v}_{\text{target}}$:
(8) $\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{K},\mathbf{v}_{\text{target}})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(\mathbf{z}_{K},\mathbf{v}_{j})/\tau)}$
The term $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\tau$ is the temperature parameter, and $B$ is the batch size, with $\{\mathbf{v}_{j}\}_{j=1}^{B}$ including the target item and negative samples from other batch items. The full training objective combines the recommendation loss with the contrastive alignment loss:
@@
where $\lambda$ is a hyperparameter that controls the contribution of the contrastive term.
@@
Table 1. Performance comparison across traditional and generative recommendation methods. Our SCoTER consistently outperforms existing baselines. Best results in bold, second-best underlined.
Baseline Methods Our Approach
Dataset Metric MF LightGCN Caser HGN Bert4Rec SASRec TIGER SCoTER Improve vs TIGER
Beauty Recall@5 0.0202 0.0228 0.0279 0.0344 0.0203 0.0387 0.0392 0.0434 10.71%
Recall@10 0.0379 0.0421 0.0456 0.0564 0.0347 0.0605 0.0594 0.0656 10.44%
NDCG@5 0.0122 0.0136 0.0172 0.0214 0.0124 0.0249 0.0257 0.0276 7.39%
NDCG@10 0.0178 0.0198 0.0229 0.0284 0.0137 0.0318 0.0321 0.0347 8.10%
Instruments Recall@5 0.0738 0.0757 0.0770 0.0813 0.0671 0.0857 0.0865 0.0908 4.97%
Recall@10 0.0967 0.1010 0.0995 0.1048 0.0822 0.1083 0.1062 0.1110 4.52%
NDCG@5 0.0473 0.0472 0.0639 0.0668 0.0560 0.0715 0.0736 0.0765 3.94%
NDCG@10 0.0547 0.0554 0.0711 0.0774 0.0608 0.0788 0.0799 0.0829 3.75%
Sports Recall@5 0.0087 0.0098 0.0116 0.0189 0.0115 0.0233 0.0233 0.0260 11.59%
Recall@10 0.0165 0.0184 0.0194 0.0313 0.0191 0.0350 0.0379 0.0406 7.12%
NDCG@5 0.0053 0.0061 0.0072 0.0120 0.0075 0.0154 0.0150 0.0161 7.33%
NDCG@10 0.0079 0.0087 0.0097 0.0159 0.0099 0.0192 0.0197 0.0209 6.09%
Yelp Recall@5 0.0220 0.0248 0.0150 0.0186 0.0186 0.0183 0.0241 0.0258 7.05%
Recall@10 0.0381 0.0403 0.0263 0.0326 0.0291 0.0296 0.0385 0.0406 5.45%
NDCG@5 0.0138 0.0156 0.0099 0.0115 0.0115 0.0116 0.0158 0.0174 10.13%
NDCG@10 0.0190 0.0207 0.0134 0.0159 0.0159 0.0152 0.0204 0.0222 8.82%
@@ 5. Experiments
RQ1: How does our proposed SCoTER framework perform against the sequential and generative recommendation models?
@@
RQ2 (What to transfer): How effective is our automated reasoning pattern discovery compared to manual, heuristic-based CoT templates?
RQ3 (How to transfer): How do structure-preserving components contribute to reasoning transfer effectiveness?
RQ4: Does integration with a backbone model synergize collaborative and reasoning signals more effectively than standalone LLM generation?

### v5

@@ 3.3.1. Structured Distillation
This stage aims to preserve the step-wise logical flow of the discovered pattern without incurring the high latency of online LLM inference. We achieve this through a structured teacher-student distillation framework.
We leverage the optimal template to guide a powerful teacher LLM (e.g., DeepSeek-R1) in generating structured reasoning chains. For each user sequence $S$ in our training corpus, the teacher model produces template-consistent reasoning $C=(c_{1},c_{2},\ldots,c_{K})$, creating training pairs $\{(S_{i},C_{i})\}_{i=1}^{N}$ where the student learns to generate structured reasoning given user sequences as input. An example synthesized instance is provided in Fig. 11.
A smaller, more efficient student model (Qwen3-8B (Yang et al., 2025a)) is fine-tuned on this synthetic dataset. This step is critical for scalability, as it allows us to generate reasoning chains for the entire user corpus at a fraction of the computational cost of the teacher model.
@@
For each generated reasoning step $c_{i,j}$, we extract a dense embedding using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025)). This process transforms textual reasoning steps into fixed-dimensional embeddings, where $\mathbf{e}_{i,j}\in\mathbb{R}^{D}$ represents the embedding for the $j$-th reasoning step of sequence $S_{i}$:
@@
All structured embedding matrices $\{\mathbf{H}_{i}\}_{i=1}^{N}$ are computed and stored offline and retrieved at inference time. This decoupling strategy eliminates the latency bottleneck of real-time generation, as the complex reasoning is reduced to a fast retrieval operation.
@@ 3.3.2. Order-Preserving Fusion
This stage integrates pre-computed step-wise matrices $\mathbf{H}_{i}$ with backbone recommendation models using a lightweight, model-agnostic fusion architecture. This online component prioritizes serving efficiency while preserving the sequential structure critical for reasoning effectiveness.
@@
To strictly enforce the logical order defined by the pattern (e.g., ensuring ”Analysis” precedes ”Recommendation”), we augment each projected embedding with learnable positional encodings:
@@
where $\mathbf{P}_{j}\in\mathbb{R}^{d_{\text{item}}}$ are position embeddings that encode each step’s role within the reasoning sequence.
We employ cross-attention to allow each sequence position to selectively attend to relevant reasoning steps. Let $\mathbf{h}_{\text{seq}}^{(i)}=[\mathbf{h}_{1}^{(i)},\mathbf{h}_{2}^{(i)},$ …, $\mathbf{h}_{T}^{(i)}]\in\mathbb{R}^{T\times d_{\text{item}}}$ denote the backbone model’s embeddings for the user sequence, and $\mathbf{Z}_{i}^{\text{pos}}=[\mathbf{z}_{i,1}^{\text{pos}};\ldots;\mathbf{z}_{i,K}^{\text{pos}}]\in\mathbb{R}^{K\times d_{\text{item}}}$ represent the projected CoT embeddings with positional encoding. In this cross-attention, the sequence embeddings serve as queries, while the reasoning steps act as both keys and values. The cross-attention mechanism computes attended reasoning representations for each sequence position:
(3) $\mathbf{A}_{i}=\text{Attention}(\mathbf{h}_{\text{seq}},\mathbf{Z}_{i}^{\text{pos}},\mathbf{Z}_{i}^{\text{pos}})$
@@
(4) $\mathbf{g}_{i}$ $=\sigma(\mathbf{W}_{g}[\mathbf{h}_{\text{seq}};\mathbf{A}_{i}]+\mathbf{b}_{g})$
(5) $\mathbf{E}_{\text{fused}}^{(i)}$ $=\text{LayerNorm}(\mathbf{g}_{i}\odot\mathbf{h}_{\text{seq}}^{(i)}+(1-\mathbf{g}_{i})\odot\mathbf{A}_{i})$
Here, $\mathbf{h}_{\text{seq}}^{(i)};\mathbf{A}_{i}]$ represents concatenation along the feature dimension, and the final layer normalization is applied to the gated output.
To align the reasoning space with the recommendation objective, we employ a contrastive learning objective. Crucially, we calculate the InfoNCE loss (Oord et al., 2018) specifically between the final reasoning step $\mathbf{z}_{i,K}^{pos}$ (which corresponds to ”Item Recommendation” in our pattern) and the target item embedding $\mathbf{v}_{\text{target}}$:
(6) $\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{i,K}^{pos},\mathbf{v}_{\text{target}})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(\mathbf{z}_{i,K}^{pos},\mathbf{v}_{j})/\tau)}$
The term $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\tau$ is the temperature parameter, and $B$ is the batch size, with $\{\mathbf{v}_{j}\}_{j=1}^{B}$ including the target item and negative samples from other batch items.
Simultaneously, we optimize the primary recommendation task. Let $p(i_{T+1}|\mathbf{E}_{\text{fused})}^{(i)}$ denote the predicted probability of the ground-truth next item $i_{T+1}$ given the fused sequence representation. The general recommendation loss is defined as:
(7) $\mathcal{L}_{\text{rec}}=-\log p(i_{T+1}|\mathbf{E}_{\text{fused}}^{(i)})$
The full training objective combines the recommendation loss with the contrastive alignment loss:
@@
where $\lambda$ controls the contrastive term.
@@
Table 1. Overall performance comparison between baselines and our proposed SCoTER. Specifically, SCoTER is applied on top of TIGER and HSTU (denoted as SCoTER-T and SCoTER-H), where “RI” indicates the relative improvement compared to the respective backbones. The best and second-best results are highlighted in bold and underlined, respectively.
Traditional Baselines TIGER Backbone HSTU Backbone
Dataset Metric LightGCN Caser HGN Bert4Rec SASRec TIGER SCoTER-T RI HSTU SCoTER-H RI
Beauty R@5 0.0228 0.0279 0.0344 0.0203 0.0387 0.0392 0.0434 10.71% 0.0567 0.0737 29.98%
R@10 0.0421 0.0456 0.0564 0.0347 0.0605 0.0594 0.0656 10.44% 0.0930 0.1187 27.63%
N@5 0.0136 0.0172 0.0214 0.0124 0.0249 0.0257 0.0276 7.39% 0.0376 0.0472 25.53%
N@10 0.0198 0.0229 0.0284 0.0137 0.0318 0.0321 0.0347 8.10% 0.0502 0.0621 23.71%
Instrument R@5 0.0757 0.0770 0.0813 0.0671 0.0857 0.0865 0.0908 4.97% 0.1009 0.1111 10.11%
R@10 0.1010 0.0995 0.1048 0.0822 0.1083 0.1062 0.1110 4.52% 0.1394 0.1573 12.84%
N@5 0.0472 0.0639 0.0668 0.0560 0.0715 0.0736 0.0765 3.94% 0.0635 0.0669 5.35%
N@10 0.0554 0.0711 0.0774 0.0608 0.0788 0.0799 0.0829 3.75% 0.0769 0.0823 7.02%
Sports R@5 0.0098 0.0116 0.0189 0.0115 0.0233 0.0233 0.0260 11.59% 0.0418 0.0584 39.71%
R@10 0.0184 0.0194 0.0313 0.0191 0.0350 0.0379 0.0406 7.12% 0.0674 0.0889 31.90%
N@5 0.0061 0.0072 0.0120 0.0075 0.0154 0.0150 0.0161 7.33% 0.0274 0.0387 41.24%
N@10 0.0087 0.0097 0.0159 0.0099 0.0192 0.0197 0.0209 6.09% 0.0361 0.0487 34.90%
Yelp R@5 0.0248 0.0150 0.0186 0.0186 0.0183 0.0241 0.0258 7.05% 0.0257 0.0356 38.52%
R@10 0.0403 0.0263 0.0326 0.0291 0.0296 0.0385 0.0406 5.45% 0.0438 0.0567 29.45%
N@5 0.0156 0.0099 0.0115 0.0115 0.0116 0.0158 0.0174 10.13% 0.0177 0.0242 36.72%
N@10 0.0207 0.0134 0.0159 0.0159 0.0152 0.0204 0.0222 8.82% 0.0245 0.0315 28.57%
@@ 4. Experiments
RQ1 (Overall Performance): How does SCoTER perform compared to existing recommendation models?
@@
RQ2 (Pattern Discovery): How effective is our automated GVM pipeline compared to manual CoT templates? Are the discovered patterns dataset-specific?
RQ3 (Structure Preservation): How do the structure-preserving components contribute to performance?
RQ4 (Robustness & Sensitivity): Is SCoTER robust across different user/item groups and hyperparameter settings?
4.1. Experimental Setup
  • HSTU (Zhai et al., 2024): A sequential model utilizing hierarchical sequential transduction units. We augment the original HSTU with semantic ID inputs and a dedicated SID decoder to enable generative retrieval (details in Appendix A.1). This enhanced version serves as a stronger baseline and backbone for SCoTER-H.
  • SCoTER: Our proposed framework. To demonstrate its universality, we implement two variants with different backbones: SCoTER-T (built upon TIGER) and SCoTER-H (built upon HSTU). Both variants enhance their respective backbones by integrating structured Chain-of-Thought reasoning, as detailed in Section 3.
Implementation Details. For traditional methods, we follow standard implementations with hyperparameters tuned on validation sets. For generative backbones, the tokenizer employs RQ-VAE (Lee et al., 2022; Zeghidour et al., 2021) for discrete semantic encoding with 4 codebooks, each containing 256 embeddings of dimension 32. Semantic inputs to RQ-VAE are derived from the embeddings of item titles and descriptions processed by Qwen3-8B-Embedding (Zhang et al., 2025). During inference, we use a beam size of 20 to balance recommendation quality and efficiency. For SCoTER, we integrate the pre-computed offline reasoning embeddings $H_{i}$ via multi-head cross-attention. We adopt distinct configurations to optimize each variant: SCoTER-T utilizes 6 cross-attention heads and is trained for 200 epochs. In contrast, SCoTER-H employs 4 cross-attention heads and is trained for 300 epochs to ensure convergence. Training uses the Adam optimizer with learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$, and a batch size of 512. The contrastive learning weight $\lambda$ is set to 0.1 for both variants. Additional hyperparameter details are provided in the Appendix A.2.
4.2. Overall Performance (RQ1)
To assess the effectiveness of our reasoning transfer framework, we compare SCoTER against strong baselines. The comprehensive results are presented in Table 1.
Consistent Improvements. SCoTER delivers consistent gains across all datasets and metrics. SCoTER-T achieves 3.75%–11.59% relative improvement over TIGER, while SCoTER-H yields larger gains of 5.35%–41.24% over the stronger HSTU backbone, achieving state-of-the-art results on most metrics.
Dataset-wise Analysis. The most substantial improvements appear on Beauty and Sports datasets, where SCoTER-H boosts Recall@5 by 29.98% and 39.71%, respectively. On Instrument, HSTU’s NDCG metrics are lower than TIGER’s, possibly due to the smaller item vocabulary favoring TIGER’s semantic tokenization. Nonetheless, SCoTER consistently improves both backbones despite these intrinsic differences. The gains are generally more pronounced for top-5 than top-10 metrics, suggesting that structured reasoning particularly benefits precision-critical scenarios.
Backbone-Agnostic Design. The consistent improvements observed across both TIGER and HSTU validate that SCoTER is backbone-agnostic, injecting effective reasoning capabilities orthogonal to the underlying base architecture.
4.3. Automated Pattern Discovery (RQ2)
To evaluate our automated GVM pipeline, we compare the discovered pattern against manual CoT templates (Two-step, Three-step, Five-step) representing increasing reasoning depth. Specifically, the Two-step template follows Rec-GPT (Yi et al., 2025). The discovered templates for each dataset are provided in Appendix F.
GVM vs. Manual Templates. As shown in Figure 3, our discovered pattern consistently outperforms all manual templates. On TIGER, SCoTER achieves 10.44% gain in Recall@10, surpassing Five-step (7.41%) by a clear margin. On HSTU, SCoTER delivers 27.63% gain versus 21.72% for Five-step. The results confirm that our data-driven approach identifies non-obvious logic pathways aligned with the data distribution, breaking through the ceiling of manual template design.
Clustering vs. Top-K Selection. We also compare our clustering-based pattern mining against a Top-K selection baseline ($K=10$) that directly selects high-scoring individual chains without clustering. As shown in Table 2, the Top-K selection leads to noticeable performance degradation. This stems from the fact that individual chain selection lacks generalizability—patterns derived from single instances fail to capture the shared reasoning structure across diverse user sequences.
Cross-Dataset Pattern Transfer. To investigate whether discovered patterns are dataset-specific, we transfer patterns across datasets. As shown in Table 3, cross-dataset transfer generally degrades performance. Notably, while the general Beauty pattern transfers reasonably well to Sports, the Sports pattern performs poorly on Beauty because it is highly domain-specific (e.g., emphasizing athletic attributes). This asymmetry confirms that effective reasoning patterns are domain-dependent, validating per-dataset discovery via GVM.
Table 2. Pattern discovery analysis on HSTU backbone.
Variant (on Beauty) R@10 N@10
Clustering (SCoTER) 0.1187 0.0621
Top-K Selection 0.1124 0.0593
Table 3. Cross-dataset pattern transfer on HSTU backbone.
Pattern Source Target R@10 N@10
Beauty Beauty 0.1187 (-) 0.0621 (-)
Sports Beauty 0.1057 ($\downarrow 10.95\%$) 0.0555 ($\downarrow 10.63\%$)
Sports Sports 0.0889 (-) 0.0487 (-)
Beauty Sports 0.0860 ($\downarrow 3.26\%$) 0.0469 ($\downarrow 3.70\%$)
Figure 3. Performance improvement relative to TIGER (top) and HSTU (bottom) backbones on the Beauty dataset. We compare manual CoT templates (Two-step, Three-step, Five-step) against our automatically discovered pattern (SCoTER).
Table 4. Ablation study of SCoTER-H on Beauty and Sports datasets. We report R@10 and N@10.
Variant Beauty Sports
R@10 N@10 R@10 N@10
SCoTER-H (Full) 0.1187 0.0621 0.0889 0.0487
w/o Position 0.1143 0.0605 0.0882 0.0479
w/o Contrastive 0.1091 0.0590 0.0842 0.0454
w/o Step-wise Embedding 0.1139 0.0607 0.0864 0.0468
w/o Step-wise (Mean) 0.1105 0.0597 0.0856 0.0462
w/o Step-wise (Last) 0.1030 0.0554 0.0827 0.0454
w/o Student Distill 0.1008 0.0537 0.0757 0.0407
HSTU (Backbone) 0.0930 0.0502 0.0674 0.0361
4.4. Effectiveness of Structure Preservation (RQ3)
To validate each component’s contribution, we conduct ablation studies on Beauty and Sports using SCoTER-H. Results are shown in Table 4. We also find that ablations on TIGER show consistent trends (see Appendix D.2).
Step-wise Fusion is Critical. Replacing step-wise attention with compressed representations causes significant degradation. Using only the last token (w/o Step-wise Last) leads to 13.23% drop in Recall@10 on Beauty, while mean pooling (w/o Step-wise Mean) results in 6.91% drop, and encoding the entire CoT as a single vector (w/o Step-wise Embedding) causes 4.04% drop. This confirms that intermediate reasoning steps contain valuable signals; discarding the logical progression reduces fine-grained matching ability.
Alignment and Position Matter. Removing contrastive alignment (w/o Contrastive) causes 8.09% drop in Recall@10 on Beauty, confirming that reasoning must be explicitly aligned with user preferences. Removing positional encoding (w/o Position) degrades Beauty by 3.71%, indicating that step order is important for logical deduction. Interestingly, the position impact on Sports is marginal, suggesting that for certain domains, semantic content dominates over strict order.
Student Distillation is Essential. Using the original Qwen3-8B (Yang et al., 2025a) without task-specific fine-tuning (w/o Student Distill) results in significantly lower performance. Specifically, on the Beauty dataset, this variant yields only an 8.39% improvement in Recall@10 over the HSTU backbone, compared to the 27.63% gain achieved with distillation. This demonstrates that task-specific adaptation is more critical than raw model capability.
4.5. Robustness and Sensitivity Analysis (RQ4)
To further analyze the robustness of our framework, we split users and items into different subgroups to gain deeper insights into SCoTER’s behavior under varying data sparsity conditions.
Setup. For users, following (Tang et al., 2025), we divide them into four equal-sized groups based on sequence length. Group 0 represents users with the sparsest interactions, while higher group numbers indicate longer sequences. For items, consistent with previous studies (Yang et al., 2023; Tang et al., 2024), we categorize them into four quartiles based on interaction frequency. Group 0 consists of the least popular items, whereas higher group numbers indicate more popular items.
Results. Figure 4 illustrates the relative improvement of SCoTER-H over HSTU on the Beauty dataset. We observe consistent performance gains across all user and item groups, demonstrating that our reasoning framework is effective regardless of data sparsity.
Figure 4. Performance comparison across different user and item groups on the Beauty dataset. Groups are divided based on interaction frequency, where Group 0 represents the sparsest scenarios and Group 3 represents the densest.
Sparse Scenarios Benefit Most. Notably, SCoTER-H achieves substantial improvements in sparse scenarios. In User Group 0, which suffers from limited historical data, the model yields a 23.90% improvement. This suggests that when collaborative signals are weak, the explicit reasoning chain helps the model deduce user intent effectively. Similarly, for Item Group 0 (tail items), we observe a remarkable 30.75% improvement. This indicates that incorporating semantic reasoning helps alleviate the popularity bias often found in generative recommenders, allowing the model to better surface relevant but less popular items.
Sweet Spot in Middle Groups. We also notice that the relative improvement peaks in middle groups (e.g., User Group 2 shows +30.24%), suggesting a sweet spot where the model effectively combines sufficient interaction history with reasoning logic to maximize performance. Even in the densest groups (Group 3), where the backbone is already strong, SCoTER maintains a significant uplift (+16.22% for users, +21.42% for items), confirming its additive value across all scenarios.
Figure 5. Impact of the contrastive learning weight $\lambda$ on Beauty and Sports datasets. The results indicate that SCoTER is relatively robust to variations in $\lambda$.
Sensitivity to Contrastive Weight $\lambda$. Figure 5 shows the impact of $\lambda$ on Beauty and Sports. Performance improves as $\lambda$ increases from 0.02 to 0.1, confirming that contrastive alignment is beneficial. The performance remains stable within 0.05–0.2 and peaks around 0.1. Further increasing $\lambda$ to 0.5 causes slight degradation, indicating that excessive alignment weight may overshadow the primary recommendation objective.
Efficiency analysis is provided in Appendix B.
Online Metric Relative Lift
GMV (Overall) +2.14%
GMV (Sparse Users) +4.10%
GMV (Dense Users) +1.49%
Negative Feedback Rate -0.24%
”Not Interested” Rate -0.25%
4.6. Online A/B Test
We validated SCoTER through deployment on the Tencent Advertising Platform. A one-week online A/B test with 5% traffic compared SCoTER against the production baseline, using Gross Merchandise Value (GMV) as the primary metric.
As reported in Table 5, SCoTER achieved a +2.14% lift in overall GMV. Stratified analysis reveals that gains are most pronounced for sparse users (+4.10%) compared to dense users (+1.49%), highlighting SCoTER’s potential to mitigate data sparsity—consistent with our offline findings. User experience also improved: negative feedback rate decreased by 0.24% and “not interested” rate by 0.25%, indicating better alignment with user preferences.
5. Conclusion
In this paper, we identify and address two challenges in applying CoT reasoning to recommendation: discovering effective reasoning patterns beyond brittle and hand-crafted heuristics, and transferring them to efficient models without collapsing their essential stepwise logic under low-latency demands. To tackle these challenges in a unified manner, we propose SCoTER, a novel framework featuring an automated GVM pipeline for pattern discovery and a structure-preserving architecture. The efficacy of this framework is validated through comprehensive experiments and real-world deployment. Comprehensive experiments demonstrate that SCoTER not only consistently outperforms state-of-the-art baselines but also achieves a 2.14% lift in production GMV. Together, these results establish SCoTER as a systematic and empirically-grounded methodology for integrating structured LLM reasoning into recommender systems.
Appendix A Experimental Setup and Hyperparameter